# PuzzleFusion++: Auto-agglomerative 3D Fracture Assembly by Denoise and Verify

### v1

@@ 4.1 Quantitative evaluations
Figure 4: Qualitative comparisons on the Breaking Bad dataset. We normalize the assembled objects for cleaner visualization so the same fragment can have different sizes in each row. For each result, we show the number of successfully assembled fragments versus the total number of fragments (i.e., the part accuracy metric). Please see the Appendix for additional results.
@@
Figure 5: Visualization of the assembly process in the first three auto-agglomerative iterations. We show two challenging examples with more than 10 fragments. The number of successfully assembled fragments increases as the system runs for more iterations.
@@ 4.2 Qualitative evaluations
Figure 4 visualizes the assembly results of PuzzleFusion++ and baselines for representative examples. The appendix provides many more results, organized by the number of fragments in each assembly without cherry-picking to provide a broad and unbiased selection. For each result, we put the number of correctly assembled fragments and the number of total fragments. We show two easy examples with less than 6 fragments in the first two rows, where Jigsaw obtains similar results to ours. The rest of the figure presents challenging examples with more than 10 fragments. As the number of fragments increases, all baselines have trouble making reasonable assembly while PuzzleFusion++ produces much more accurate and reasonable results. Specifically, the potential ambiguity in local geometric features could deteriorate the performance of Jigsaw when there are too many small fragments, while our approach stays robust by exploiting high-level geometric reasoning in the auto-agglomerative process. Please refer to Appendix B for direct comparisons between Jigsaw and our PuzzleFusion++, and the supplementary video for the animations.
4.3 Ablation study
@@
Iterations of the auto-agglomerative process. Table 2 presents the ablation study on the number of auto-agglomerative iterations (#$ite$): The second row (#$ite$=1) demonstrates that our denoiser alone performs better than the other baselines (Table 1). DiffAssemble[28] also uses diffusion models but performs much worse, indicating the effectiveness of our feature embeddings (§3.2) and system-level designs (Figure 3). The rest of the table shows that our performance consistently improves with more auto-agglomerative iterations (i.e., iterating the denoiser and the verifier). The iterations are particularly important for complex assemblies with many fragments, as shown by two hard examples in Figure 5. The verifier finds a few small fragments correctly aligned in the early iterations, which simplifies the problem for the denoiser and helps complete the assembly in the later iterations.
@@
Sampling steps of the denoiser. Table 3 shows how performance varies with different numbers of sampling steps when only the denoiser is used during testing. The optimal performance is achieved with 20 sampling steps, yet performance remains robust even with fewer steps. Accelerated sampling in diffusion models[33] is an active area of research, leaving good potential for our approach also to achieve a boost in speed.
@@ 4.4 Failure cases and limitations
Figure 6 shows two failure modes due to 1) local geometric ambiguity and 2) small fracture surfaces.
$\bullet$ Local geometric ambiguity (left four examples) result in the misplacement of fragments such as the red and yellow fragments on a mug handle in the first column, the green and yellow bowl fragments in the second column, and an upside-down bottle in the fourth column. The third column presents an exceptionally difficult scenario with 20 fragments. As the number of fragments increases, the likelihood of encountering fragments with similar local geometries also rises. Nonetheless, our method uses global shape priors to assemble all fragments into a coherent global shape.
@@ 5 Conclusion
Social impacts. PuzzleFusion++ brings positive social impacts. In archaeology, it can aid in reconstructing ancient artifacts and preserving cultural heritage. In medicine, it can enhance preoperative planning by accurately reconstructing fractured bones from CT scans. Forensic science can benefit from more precise reconstruction of evidence from accident scenes, aiding in criminal investigations. Additionally, in scientific research, PuzzleFusion++ can improve the structural analysis of proteins, potentially accelerating drug discovery. On the negative side, automation of such complex tasks, traditionally involving human cognitive and manual skills, could lead to the displacement of jobs.
Appendix
The appendix provides the remaining system details and additional experimental results. Please also refer to the qualitative results on our project page for a detailed demonstration of PuzzleFusion++’s auto-agglomerative assembly process.
@@ A.1 Details of the Fragment Autoencoder
We employ PointNet++ for self-supervised pre-training of the autoencoder, using a Single-Scale Grouping PointNet++ encoder $E$ to aggregate features. Each fragment is represented by a point cloud with 1000 uniformly sampled points. The point cloud is normalized so the center of mass is at the origin and the longest dimension of the axis-aligned bounding box becomes a unit length. The point cloud is initially encoded into 25 distinct point latent vectors representing local features with dimension 64. Each point latent $z_{p}$ is associated with the local center coordinates $c_{p}$.
@@
A.2 Details of noise scheduler
We have presented our noise scheduler in Figure 3 of the main paper. Similar to humans solving jigsaw puzzles, accurately aligning local fracture surfaces is usually more challenging than knowing the rough location of a fragment. Therefore, we allocate more denoising budgets to getting precise alignments than moving fragments to the rough locations by designing the piecewise function. Figure 7 shows the comparison between the commonly used linear scheduler, cosine scheduler, and our customized scheduler.
Figure 7: Our customized noise scheduler.
@@ Appendix B Additional Experimental Results and Analyses
We provide additional qualitative results based on the number of fragments: Figure 8 for 2-5 fragments, Figure 9 for 6-10 fragments, Figure 10 for 11-15 fragments, and Figure 11 for 16-20 fragments. In the main paper, we have demonstrated that Jigsaw is the primary competing method to our method, so we exclude other baselines to save space.
@@
Figure 8: Additional qualitative results for 2 to 5 fragments.
Figure 9: Additional qualitative results for 6 to 10 fragments.
Figure 10: Additional qualitative results for 11 to 15 fragments.
Figure 11: Additional qualitative results for 16 to 20 fragments.

### v2

@@ 4.1 Quantitative evaluations
Figure 4: Qualitative comparisons on the Breaking Bad dataset (mesh used only for visualization). We transform the assembled objects to same coordinate system and normalize their sizes for clearer visualization. For each result, we show the number of successfully assembled fragments versus the total number of fragments (i.e., the part accuracy metric). Please see the Appendix for additional results.
@@
Table 4: Ablation study on the core design of the denoiser.
Scheduler Autoencoder Anchor RMSE (Rot.) RMSE (Trans.) PA $\times$ 48.2 11.9 57.5 $\times$ 53.5 10.5 56.9 $\times$ 47.4 10.7 61.1 40.8 9.06 67.3
Table 5: Ablation study on different types of verifier.
Method Verifier Type RMSE (Rot.) RMSE (Trans.) PA Ours (#$ite$=1) None 40.8 9.06 67.3 Ours (#$ite$=6) Jigsaw 38.1 8.04 70.6 Ours (#$ite$=6) GT (Upper Bound) 34.0 5.87 82.9
Table 6: Ablation on verifier’s performance
Threshold Accuracy (%) Precision (%) Recall (%) 0.5 87.99 58.64 73.51 0.9 90.77 87.88 45.95
Table 7: Random initializations of the anchor fragment.
RMSE(R) RMSE(T) PA CD $40.86\pm 0.37$ $9.03\pm 0.18$ $67.49\pm 0.51$ $6.69\pm 0.62$
@@ 4.2 Qualitative evaluations
Figure 4 visualizes the assembly results of PuzzleFusion++ and baselines for representative examples. The §B provides many more results, organized by the number of fragments in each assembly without cherry-picking to provide a broad and unbiased selection. For each result, we put the number of correctly assembled fragments and the number of total fragments. We show two easy examples with less than 6 fragments in the first two rows, where Jigsaw obtains similar results to ours. The rest of the figure presents challenging examples with more than 10 fragments. As the number of fragments increases, all baselines have trouble making reasonable assembly while PuzzleFusion++ produces much more accurate and reasonable results. Specifically, the potential ambiguity in local geometric features could deteriorate the performance of Jigsaw when there are too many small fragments, while our approach stays robust by exploiting high-level geometric reasoning in the auto-agglomerative process. Please refer to §B for direct comparisons between Jigsaw and our PuzzleFusion++, and the supplementary video for the animations.
4.3 Ablation studies and analyses
Iterations of the auto-agglomerative process. Table 7 presents the ablation study on the number of auto-agglomerative iterations (#$ite$): The second row (#$ite$=1) demonstrates that our denoiser alone performs better than the other baselines (Table 1).
Figure 5: Visualization of the assembly process in the first three auto-agglomerative iterations (mesh used only for visualization). We show two challenging examples with more than 9 fragments. The number of successfully assembled fragments increases as the system runs for more iterations.
@@ 4.3 Ablation studies and analyses
The rest of the table shows that our performance consistently improves with more auto-agglomerative iterations (i.e., iterating the denoiser and the verifier). The iterations are particularly important for complex assemblies with many fragments, as shown by two hard examples in Figure 10. The verifier finds a few small fragments correctly aligned in the early iterations, which simplifies the problem for the denoiser and helps complete the assembly in the later iterations.
@@
Denoiser. Table 7 shows how performance varies with different numbers of sampling steps when only the denoiser is used during testing. The optimal performance is achieved with 20 sampling steps, yet performance remains robust even with fewer steps.
DiffAssemble (Scarpellini et al., 2024) was initially designed to solve 2D jigsaw puzzles using a diffusion model, and it directly extends the 2D architecture to tackle the 3D variant without proper designs for 3D. While achieving SOTA performance on the 2D jigsaw puzzle, it fails to produce reasonable results on the 3D task – our denoiser-only result outperforms it by a huge margin (67.3% vs. 27.5%). Table 7 shows the ablation of our key denoiser designs: 1) pre-training the autoencoder to learn better local geometric features for fractures, 2) introducing an anchor fragment to resolve the ambiguity in 3D assembly, and 3) adjusting the sampling scheduler to balance the rough and fine alignments. These three changes bring performance closer to the baseline method, Jigsaw. Additionally, our research contribution (auto-agglomerative formulation) provides a significant advantage over Jigsaw in addition to these improvements.
Verifier. Table 7 shows the performance of the verifier with different thresholds. In practice, we use the threshold of 0.9 for the transformer’s output to be classified as correct alignment. This high threshold ensures that the selected pairs have high confidence, which is crucial since our framework does not backtrack errors. The verifier’s accuracy is 90.77%, which is high because all pairs are input into the transformer, and most pairs are trivially not matched (i.e., they do not have any matching points). The lower recall (45.95%) indicates the model misses many true matches.
The suboptimal accuracy of the verifier limits the gain brought by increasing the number of auto-agglomerative iterations. To understand the upper bound of our method, we experiment with a “ground-truth verifier”, which always makes the correct merging of pieces. Note that this verifier does not leak ground-truth pose information to the denoiser but rather provides correct high-level piece-merging guidance. Interestingly, Table 7 shows that this leads to significantly better performance than the verifier using Jigsaw matchings. A more robust verifier model, as well as a potential backtracking strategy, are promising directions for future work.
Random initialization of the anchor fragment. Table 7 presents the results of 10 model inferences with different random initializations of the anchor fragment and calculates the mean and standard deviation of the metrics. Different anchor fragment initialization does not affect the quality of the assembly results. In addition, Figure 6 shows one example assembly process with different anchor initialization. The anchor fragment is randomly rotated and moved to the origin during both training and testing. As explained in §3.1, it helps set up a fixed reference coordinate frame for the assembly process, resolving the inherent ambiguity in 3D rigid transformation.
Figure 6: Assembled results of a single example with different anchor initializations from timestep T to 0. The red circle marks the anchor fragment, with the object assembled around it.
@@ 4.4 Failure cases and limitations
Figure 7 shows two failure modes due to 1) local geometric ambiguity and 2) small fracture surfaces. We discuss the two challenges and the limitations of our approach below:
$\bullet$ Local geometric ambiguity (left four examples) result in the misplacement of fragments such as the red and yellow fragments on a mug handle in the first column. The third column presents an exceptionally difficult scenario with 20 fragments. As the number of fragments increases, the likelihood of encountering fragments with similar local geometries also rises. Nonetheless, our method uses global shape priors to assemble all fragments into a coherent global shape.
@@ 5 Conclusion
Acknowledgments
This research is partially supported by NSERC Discovery Grants, NSERC Alliance Grants, and John R. Evans Leaders Fund (JELF). We thank the Digital Research Alliance of Canada and BC DRI Group for providing computational resources.
Appendix: PuzzleFusion++: Auto-agglomerative 3D
Fracture Assembly by Denoise and Verify
The appendix provides the remaining system details and additional experimental results. Please also refer to our project page https://puzzlefusion-plusplus.github.io for a detailed video demonstration of PuzzleFusion++’s auto-agglomerative assembly process.
@@ A.1 Details of the Fragment Autoencoder
Figure 8 shows the reconstruction results of our autoencoder. We employ PointNet++ for self-supervised pre-training of the autoencoder, using a Single-Scale Grouping PointNet++ encoder $E$ to aggregate features. Each fragment is represented by a point cloud with 1,000 uniformly sampled points. The point cloud is normalized so the center of mass is at the origin and the longest dimension of the axis-aligned bounding box becomes a unit length. The point cloud is initially encoded into 25 distinct point latent vectors representing local features with dimension 64. Each point latent $z_{p}$ is associated with the local center coordinates $c_{p}$.
@@
Figure 8: Reconstruction results of our point clouds VQ-VAE. The comparison highlights the quality differences between the original point cloud and the reconstructed point cloud generated by the VQ-VAE. In the reconstructed point cloud, the same color represents points reconstructed by the same latent. Each latent is responsible for reconstructing only a local area. This figure demonstrates that the key features of fracture fragments can be preserved in the latent codes.
A.2 Details of the noise scheduler
We have presented our noise scheduler in Figure 3 of the main paper. Similar to humans solving jigsaw puzzles, accurately aligning local fracture surfaces is usually more challenging than knowing the rough location of a fragment. Therefore, we allocate more denoising budgets to getting precise alignments than moving fragments to the rough locations by designing the piecewise function. Figure 9 plots the noise scheduling curves for the linear scheduler, the cosine scheduler, and our customized scheduler. Figure 10 provides the visualization of the denoising process for a vase model. Figure 11 and Table 8 provide the qualitative and quantitative comparisons of the three noise schedulers, respectively.
Our noise scheduler is used for both training and testing, as we follow the vanilla DDPM formulation rather than EDM (Karras et al., 2022). With our tailored noise scheduler for the 3D shape assembly task, the denoising process allocates more steps to refining precise local adjustments rather than finding the rough global location (at test time). As for training, this scheduler indeed makes the process more efficient, as fewer training iterations are spent on "less important" timesteps. If the default (linear or cosine) scheduler were used for training while our scheduler was applied during testing, similar results might still be achieved but would require more training iterations.
To illustrate the advantages of our scheduler during the testing stage, we compare it to the linear and cosine scheduler. The linear scheduler uses most of its steps (T=1000 to 150) for rough localization, while the cosine scheduler allocates more denoising steps in the final adjustment phase, outperforming the linear scheduler by a clear margin .
Building on this idea, our scheduler dedicates an even larger portion of the denoising steps to the final adjustment, further improving results. The top three rows of Figure 11 show simpler cases of objects comprising at most 5 fragments. While all the schedulers achieve 100% part accuracy, gaps between fragments are visible for the linear or the cosine schedulers. Our precise alignments may have minimal effects on the standard metrics but significantly enhance the quality of the final assembly.
Figure 9: Comparing different noise schedulers. Our customized scheduler adds large noise near the end of the diffusion steps. During denoising, we spend more iterations at low noise levels, figuring out precise final alignments.
Figure 10: Comparing different noise schedulers. The figure shows the denoising process for a vase model for the three noise schedulers shown in Figure 9. The linear scheduler uses most of the denoising steps (T=1000 to T=150) to find a rough location, leaving a few steps for alignment. The cosine scheduler improves efficiency but still wastes steps on rough localization. In contrast, our scheduler quickly identifies the rough location and dedicates more steps to precise alignment.
Table 8: Comparing different noise schedulers. Quantitative evaluations of the three noise schedulers shown in Figure 9. Our customized scheduler spends more denoising steps at low noise levels, achieving the best numbers in all the metrics.
Scheduler RMSE (Rot.) $\downarrow$ RMSE (Trans.) $\downarrow$ PA $\uparrow$ CD $\downarrow$ Linear 48.2 11.9 57.5 9.21 Cosine 43.4 9.24 64.4 7.10 Ours 40.8 9.06 67.3 6.45
Figure 11: Comparing different noise schedulers. The figure shows the final assembly results of the three noise schedulers shown in Figure 9. The first three rows highlight simple cases where our scheduler achieves precise alignment, surpassing others with visible gaps. The last three rows showcase challenging cases, demonstrating the superior performance of our scheduler.
@@ Appendix B Additional Experimental Results and Analyses
B.1 More analysis on Everyday object subset
More qualitative results. We provide additional qualitative results based on the number of fragments: Figure 12 for 2-5 fragments, Figure 13 for 6-10 fragments, Figure 14 for 11-15 fragments, and Figure 15 for 16-20 fragments. In the main paper, we have demonstrated that Jigsaw is the primary competing method to our method, so we exclude other baselines to save space.
@@ B.1 More analysis on Everyday object subset
Figure 12: More final assembly results for objects comprising 2 to 5 fragments.
Figure 13: More final assembly results for objects comprising 6 to 10 fragments.
Figure 14: More final assembly results for objects comprising 11 to 15 fragments.
Figure 15: More final assembly results for objects comprising 16 to 20 fragments.
More quantitative analysis. Figure 16 shows the distribution of the part accuracy metric with respect to the number of fragments compared with Jigsaw. Our performance drops when the number of fragments increases, as there could be many small pieces and, leading to much higher difficulty. However, compared to Jigsaw, our performance is still significantly better for a large number of fragments.
Figure 16: Part accuracy grouped by the number of fragments: Results are tested on the everyday object subset. Shaded areas show part accuracy distribution of our method and Jigsaw. Solid lines indicate the average. Our method presents clear superiority for complex objects with more fragments.
B.2 More analysis on complicated object
The Everyday subset contains relatively simple objects, such as bottles, glasses and wine. To explore the model’s ability to handle more complex shapes, we focused on the Artifact subset, which features challenging archaeological objects with intricate structures. To evaluate this, we fine-tuned the model, initially trained on the Everyday subset, on the Artifact subset.
Table 9 presents the quantitative evaluation on the Artifact subset, comparing our method with baseline approaches: Global, LSTM, and DGL. Our method achieves significantly better performance than these methods. Following this, Figure 17 showcases the qualitative results, further demonstrating the effectiveness of our approach on the challenging object.
Table 9: The Artifact subset of the Breaking Bad dataset contains objects with complicated shapes. We trained the models on the Everyday subset, fine-tuned on the Artifact subset, then evaluated on the Artifact subset in this table.
Method RMSE (Rot.) $\downarrow$ RMSE (Trans.) $\downarrow$ PA $\uparrow$ CD $\downarrow$ Global 83.8 16.6 19.0 13.3 LSTM 84.6 16.8 21.5 11.7 DGL 81.7 16.6 17.3 19.4 Ours (#$ite$=1) 41.1 9.87 62.6 9.29
Figure 17: Final assembly results of 12 complex objects from the Artifact subset, including intricate archaeological artifacts such as statues, animal figures, and ornamental designs.
B.3 Compared with other recent baselines
3D Geometric Shape Assembly via Efficient Point Cloud Matching. PMTR (Lee et al., 2024) evaluated their method on a simplified version of the data in the Breaking Bad dataset, called the volume-constrained version. This version filters out fragments below a minimum volume threshold, effectively removing very small fragments and making the dataset less challenging. To ensure a fair evaluation, we trained our method on the same volume-constrained Everyday subset and reported the results in Table 10. Our method outperforms PMTR across all metrics, except for the rotation error (RMSE(R)), where it lags by a small margin of approximately 1 degree.
Table 10: PMTR Lee et al. (2024) used a simplified version of the Breaking Bad dataset, called the volume-constrained version. The version filters out fragments below a minimum volume threshold, making the version easier than the standard set. For fair comparison, the table compares our system and PMTR on the volume-constrained version of the Everyday subset.
Method RMSE (Rot.) $\downarrow$ RMSE (Trans.) $\downarrow$ PA $\uparrow$ CD $\downarrow$ PMTR(Lee et al., 2024) 31.57 9.95 70.6 5.56 Ours 32.70 5.41 78.9 3.01
Multi-Fragment Assembly Using Proxy-Level Hybrid Transformer. PHFormer (Cui et al., 2024) employs a hybrid attention module to model the relationships between fragments. We compared our performance with theirs in the Table 11.
Table 11: Comparison of our method and PHFormer performance on the Everyday subset
Method RMSE (Rot.) $\downarrow$ RMSE (Trans.) $\downarrow$ PA $\uparrow$ CD $\downarrow$
PHFormer 26.1 9.3 50.7 9.6
Ours 38.1 8.04 70.6 6.02
Our method outperforms PHFormer in three metrics but has a higher RMSE in rotation.
Fracture Assembly with Segmentation and Iterative Registration. FRASIER (Kim et al., 2024) reassembles fractured objects using fracture surface segmentation and iterative registration. It uses GeoTransformer (Qin et al., 2022) to match points between fragment pairs and samples 50k points per fragment. This high point density allows FRASIER to capture detailed fracture surfaces but does not align with the standard 1k-point setup used by baseline methods.
To infer FRASIER’s performance under the 1k-point setting, we refer to GeoTransformer’s reported performance in Jigsaw (Lu et al., 2023) (Appendix E.3). With 1k points, GeoTransformer produces poor results (RMSE (Rot.) = $84.8^{\circ}$, RMSE (Trans.) = $14.3\times 10^{-2}$, PA = 3.1%). Since FRASIER depends on GeoTransformer for registration, it likely cannot deliver reasonable results under the 1k-point-per-fragment setting.
B.4 Analysis on high rotation errors
Table 7 shows the upper bound of our method, which leverages the ground truth verifier. However, the rotation error remains high (34-degree error). We perform further analysis compared our method and Jigsaw using the results from Table 1. Table 12 shows the improvement margin for rotation is relatively modest (+9.93%), whereas other metrics show significant gains exceeding 20%.
Table 12: Improvement margins (Delta) calculated as the relative performance difference between our method and Jigsaw. Positive values indicate that our method outperforms the baseline.
RMSE (Rot.) $\downarrow$ RMSE (Trans.) $\downarrow$ PA $\uparrow$ CD $\downarrow$
Delta +9.93% +24.86% +23.21% +54.66%
In addition, the failure cases illustrated in Figure 7 are also highly related to the high rotation error:
  • Local geometric ambiguity: Similar geometry across fragments makes it difficult to determine precise rotations.
  • Small fracture surfaces: Tiny pieces often lack distinct surface features, leading to 180-degree rotation errors.
All these results demonstrate that our diffusion-based method does not handle rotation as effectively as traditional optimization-based approaches. We provide some thoughts below:
The other three metrics heavily rely on an accurate translation/placement of fragment pieces, where the diffusion-based approach excels by learning global shape priors together with local alignments. On the contrary, the RMSE of rotation mainly examines the accuracy of fine-grained alignments, which relies more on accurate local shape matching. It seems that methods like Jigsaw can produce better rotation-level alignments by conducting direct optimization based on the local surface matching results, while our diffusion-based approach allocates most of the learning capacity for global shapes (correct translations). We believe that a potential direction for improving the rotation-level accuracy can be adding an additional stage that focuses on refining the rotation parameters.